For numbers 13 and 14, the length of a rectangular basketball court
is 5 meters more than its width. The area of the court is 750 square
meters and the width of the court is x meters.
13. Which equation can be used to find the width of the basketball
court?
Ⓐ x² + 5x + 750 = 0
Ⓑ x² − 5x + 750 = 0
Ⓒ x² + 5x − 750 = 0
Ⓓ x² − 5x − 750 = 0
14. What is the width of the rectangular court in meters?

Answer by Theo(13342)

length = x + 5
width = x

area = length * width = x * (x + 5) = x^2 + 5x.
since area = 750, you get x^2 + 5x = 750
subtract 750 from both sides of that equation to get x^2 + 5x - 750 = 0
that would be selection C.

factor that equation to get (x + 30) * (x - 25) = 0
solve for x to get x = -30 or x = 25
x = 25 because x can't be negative.

your solution is the width of the court is x = 25 meters.